Find the inverse function of (x+1)/x

1
Expert's answer
2022-05-03T07:50:10-0400

The domain of the function "f(x)=\frac{x+1}{x}" consists of all real numbers except of . I.e., the domain is: "(-\infty,0)\cup(0,+\infty)". In order to find the inverse, we need to solve the equation: "z=\frac{x+1}{x}". We get: "zx=x+1". We receive: "x=\frac{1}{z-1}". The domain of the inverse function is: "(-\infty,1)\cup(1,+\infty)"

Answer: the inverse function is: "\frac{1}{z-1}".